# Block-Diagonal Orthogonal Relation and Matrix Entity for Knowledge Graph Embedding

### v1

@@ 5.1 Experiment Setup
We evaluate our proposed method on two KG datasets, including WN18RR Dettmers et al. (2018), FB15K-237 Toutanova and Chen (2015). The details of these datasets are shown in Table 1.
@@
We compare our new model with state-of-the-art (SOTA) methods, namely TransE Bordes et al. (2013), DistMult Yang et al. (2014), ComplEx Trouillon et al. (2016), ConvE Dettmers et al. (2018), RotatE Sun et al. (2019), and QuatE Zhang et al. (2019). In addition to these five and our OrthogonalE, we introduce comparative models utilizing the Gram-Schmidt process for generating orthogonal matrices and SGD for joint relation-entity training. OrthogonalE further differentiates by employing orthogonal matrices of varying sizes to discuss performance nuances.
@@
The key hyperparameters in our implementation include the learning rate for RiemannianAdam Kochurov et al. (2020) and Adagrad Duchi et al. (2011), negative sample size, and batch size. To determine the optimal hyperparameters, we performed a grid search using the validation data.
Figure 4: MRR accuracy comparison of OrthogonalE models with different small orthogonal matrices across varying entity dimensions (n1n*1, we set $m=1$) on WN18RR.
Figure 5: MRR accuracy comparison of OrthogonalE models with different small orthogonal matrices across varying entity dimensions (n1n*1, we set $m=1$) on FB15K-237.
Figure 6: MRR accuracy comparison of OrthogonalE($\times2$) and Gram-Schmidt($\times2$) models across varying entity dimensions ($m$) with fixed relation matrix ($\times40$) on WN18RR.
Figure 7: MRR accuracy comparison of RotatE and OrthogonalE($\times2$) models across varying entity dimensions (nmn*m) on WN18RR.
Figure 8: MRR accuracy comparison of QuatE and OrthogonalE($\times3$) models across varying entity dimensions (nmn*m) on WN18RR.
Figure 9: MRR accuracy comparison of RotatE and OrthogonalE($\times2$) models across varying entity dimensions (nmn*m) on FB15K-237.
@@ 5.2 Results
We first analyzed the overall accuracy for all baseline models and OrthogonalE, then separately examined the impacts of Block-diagonal Orthogonal matrices, Riemannian Optimization for relations, and Entity matrices on the model from various experimental results. These results provided evidence supporting OrthogonalE’s generality and flexibility.
Overall Accuracy
@@
Table 2 presents link prediction accuracies for the WN18RR and FB15K-237 datasets. The OrthogonalE model surpasses all state-of-the-art models, highlighting its superior representational ability by employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Moreover, the OrthogonalE model with $\times2$ and $\times3$ configurations yields significantly better performance than corresponding sizes of the Gram-Schmidt method, and notably exceeds RotatE and QuatE, respectively, showcasing the enhanced efficacy of the KGE model.
Block-diagonal Orthogonal matrices
@@
Figure 4 and 5 show MRR accuracy comparison of OrthogonalE models with different block-diagonal orthogonal matrices across varying entity dimensions (n1n*1, we set $m=1$) on WN18RR and FB15K-237, repectively.
An initial analysis of the datasets reveals that WN18RR contains 40,943 entities but only 11 relations, averaging 3,722 entities per relation. Conversely, FB15K-237 has 14,541 entities and 237 relations, averaging only 61 entities per relation. This suggests that WN18RR, in comparison to FB15K-237, requires a more complex and effective representation capability.
@@
Our experimental findings align with the dataset analysis. For WN18RR, as shown in Figure 4, there is little difference in performance between $\times3$, $\times4$, $\times5$, and $\times10$ blocks, but these configurations significantly outperform the $\times2$ blocks. In contrast, for FB15K-237 depicted in Figure 5, the results across block-diagonal sizes, including $\times2$, $\times3$, $\times4$, and $\times10$ are consistent. This suggests that $\times2$ blocks suffice for representing FB15K-237’s relations, while WN18RR requires higher-dimensional, more complex blocks, such as $\times3$ or $\times4$, for adequate representation. These results illustrate that the OrthogonalE model is general, which can adapt to datasets of various complexities by adjusting the dimension $d$ of the block-diagonal matrices.
Riemannian Optimization for relations
@@
Figure 6 compares MRR accuracies of OrthogonalE ($\times2$) and Gram-Schmidt ($\times2$) across entity dimensions ($m$) with a constant relation matrix ($\times40$) on WN18RR, assessing the efficacy of orthogonal optimization beyond the Gram-Schmidt method for block-diagonal orthogonal matrices. The result demonstrates that OrthogonalE’s Riemannian optimization significantly exceeds Gram-Schmidt, underscoring its necessity.
Entity matrix
@@
Figures 7 and 8 compare MRR accuracies of RotatE with OrthogonalE ($\times2$) and QuatE with OrthogonalE ($\times3$) over different entity dimensions nmn*m on WN18RR, then Figure 9 shows comparison of RotatE and OrthogonalE ($\times2$) in FB15-237. In OrthogonalE, we maintained a constant entity dimension (nmn*m) while varying $m$ to assess the impact of entity shape. Figures 7 and 9 show that models with mm=1, 2, or 3(WN18RR) perform similarly and outperform $m=10$; notably, even when the number of relation parameters is just 1/10th of RotatE’s, the performance is significantly better at all dimensions. Figure 8 mirrors these findings, with mm=1 and 2 outperform mm=10, and mm=10 again surpassing QuatE, which has 7.5 times more parameters. These results demonstrate OrthogonalE’s efficacy in saving relation parameters while outperforming RotatE and QuatE, highlighting our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
@@
Furthermore, Figure 6 result (with $m=7$ yielding MRR=0.483) suggests that a relation matrix of $\times40$ (20 parameters), compared to a dimension of 500 (250 parameters), can achieve comparably high performance, thus demonstrating that entity matrix method significantly reduces the need for excessive relation parameters to yield effective outcomes.
@@ 6 Conclusion
In this study, we propose the OrthogonalE model to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. The experimental results demonstrate that our new KGE model, OrthogonalE, is both general and flexible, significantly outperforming state-of-the-art KGE models while substantially reducing the number of relation parameters.

### v2

@@ 5.1 Experiment Setup
We evaluate our proposed method on two KG datasets, including WN18RR Dettmers et al. (2018) (license: Apache 2.0), FB15K-237 Toutanova and Chen (2015) (license: CC-BY-4.0). The details of these datasets are shown in Table 4. More detail refers to A.1.
@@
We compare our new model with classic methods, including TransE Bordes et al. (2013), DistMult Yang et al. (2014), ConvE Dettmers et al. (2018), and rotation-based KGE methods such as ComplEx Trouillon et al. (2016), RotatE Sun et al. (2019), QuatE Zhang et al. (2019), HopfE Bastos et al. (2021), and DensE Lu et al. (2022). In addition to these methods and our OrthogonalE$(d\times d)$, we introduce comparative models Gram-Schmidt$(d\times d)$ utilizing the Gram-Schmidt process for generating orthogonal matrices and SGD for joint relation-entity training. OrthogonalE further differentiates by employing orthogonal matrices of varying sizes to discuss performance nuances.
@@
The key hyperparameters in our implementation include the learning rate for RiemannianAdam Kochurov et al. (2020) and Adagrad Duchi et al. (2011), negative sample size, and batch size. To determine the optimal hyperparameters, we performed a grid search using the validation data. More detail refers to A.1.
Figure 4: MRR accuracy comparison of OrthogonalE models with different small orthogonal matrices across varying entity dimensions ($n*1$, we set $m=1$) on WN18RR and FB15K-237.
@@ 5.2 Results
We first analyzed the overall accuracy for all baseline models and OrthogonalE, then separately examined the impacts of Block-diagonal Orthogonal matrices, Riemannian Optimization for relations, and Entity matrices on the model from various experimental results. Finally, we utilize several relation histograms to verify our model can capture these relation patterns.
5.2.1 Overall Accuracy
@@
Table 3 presents link prediction accuracies for the WN18RR and FB15K-237 datasets. The OrthogonalE model demonstrates superior performance in the WN18RR dataset and achieves results on the FB15K-237 dataset that are only marginally lower than those of HopfE Bastos et al. (2021), outperforming all other compared models, highlighting its superior representational ability by employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Moreover, the OrthogonalE model with 2$\times$2 and 3$\times$3 configurations yields significantly better performance than the corresponding sizes of the Gram-Schmidt method, and notably exceeds RotatE and QuatE, respectively, showcasing the enhanced efficacy of the KGE model. Finally, since the WN18RR and FB15K-237 datasets are relatively small, the performance differences among OrthogonalE models with ($2\times 2$), ($3\times 3$), ($4\times 4$), and ($10\times 10$) are not significant when using sufficient dimensions (WN18RR: 500, FB15K-237: 1000). We will discuss the performance at different dimensions in detail in section 5.2.2.
5.2.2 Block-diagonal Orthogonal matrices
@@
Fig. 4 shows MRR accuracy comparison of OrthogonalE models with different block-diagonal orthogonal matrices across varying entity dimensions ($n*1$, we set $m=1$) on WN18RR and FB15K-237.
An initial dataset analysis reveals WN18RR has 40,943 entities with just 11 relations (about 3,722 entities per relation), while FB15K-237 includes 14,541 entities and 237 relations (around 61 entities per relation). This implies WN18RR requires a more sophisticated representation capability compared to FB15K-237.
@@
Our results (Fig. 4) confirm our dataset analysis. For WN18RR, performance is similar for block sizes from 3$\times$3 to 10$\times$10, all outperforming 2$\times$2 blocks, showcasing 2$\times$2 blocks are not enough for its relation representation. However, for FB15K-237, performance is stable across all block sizes, indicating 2$\times$2 blocks are enough for its relations representation. These results show WN18RR requires more complex blocks for adequate representation, and illustrate that the OrthogonalE model is general, which can adapt to datasets of various complexities by adjusting the dimension $d$ of the block-diagonal matrices.
Figure 5: MRR accuracy comparison of OrthogonalE(2$\times$2) and Gram-Schmidt(2$\times$2) models across varying entity dimensions ($m$) with fixed relation matrix (40$\times$40) on WN18RR.
Figure 6: MRR accuracy comparison of RotatE and OrthogonalE(2$\times$2) models across varying entity dimensions ($n*m$) on WN18RR.
Figure 7: MRR accuracy comparison of QuatE and OrthogonalE(3$\times$3) models across varying entity dimensions ($n*m$) on WN18RR.
Figure 8: MRR accuracy comparison of RotatE and OrthogonalE(2$\times$2) models across varying entity dimensions ($n*m$) on FB15K-237.
Figure 9: Histograms of relation embeddings for different relation patterns, where $\mathbf{e}_{R}^{1}$ represents _similar_to, $\mathbf{e}_{R}^{2}$ represents _member_of_domain_region, $\mathbf{e}_{R}^{3}$ represents /film/film/genre, $\mathbf{e}_{R}^{4}$ represents /media_common/netflix_genre/titles, $\mathbf{e}_{R}^{5}$ represents /location/administrative_division/country, $\mathbf{e}_{R}^{6}$ represents /location/hud_county_place/place, and $\mathbf{e}_{R}^{7}$ represents /base/aareas/schema/administrative_area/capital. From the WN18RR dataset, we select $\mathbf{e}_{R}^{1}$ and $\mathbf{e}_{R}^{2}$ to represent Symmetry and Antisymmetry, respectively, and obtain their relation embeddings using the OrthogonalE(3$\times$3) model with $n$=500 and $m$=1. Similarly, from the FB15K-237 dataset, we select $\mathbf{e}_{R}^{3}$, $\mathbf{e}_{R}^{4}$, and $\mathbf{e}_{R}^{5}$, $\mathbf{e}_{R}^{6}$, $\mathbf{e}_{R}^{7}$ as representations for Inversion and Composition, respectively, and acquire their relation embeddings under the OrthogonalE(2$\times$2) model with $n$=1000 and $m$=1.
5.2.3 Riemannian Optimization for relations
@@
Fig. 5 compares MRR accuracies of OrthogonalE (2$\times$2) and Gram-Schmidt (2$\times$2) across entity dimensions ($m$) with a constant relation matrix (40$\times$40) on WN18RR, assessing the efficacy of orthogonal optimization beyond the Gram-Schmidt method for block-diagonal orthogonal matrices. The result demonstrates that OrthogonalE’s Riemannian optimization significantly exceeds Gram-Schmidt, underscoring its necessity.
5.2.4 Entity matrix
@@
In OrthogonalE, we maintained a constant entity dimension ($n*m$) while varying $m$ to assess the impact of entity shape. Fig. 6 compare MRR accuracies of RotatE with OrthogonalE (2$\times$2) over different fixed entity dimensions $n*m$ on WN18RR. OrthogonalE models with $m=1$, $2$, or $3$ perform similarly and better than $m=10$, and all significantly outperform RotatE across dimensions. Notably, their relation parameter is $1/m$ of RotatE’s, which shown in Table. 2. These results demonstrate OrthogonalE’s efficacy in saving relation parameters while outperforming RotatE, highlighting our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
Besides the comparison of RotatE and OrthogonalE(2$\times$2), Fig. 7 compare MRR accuracies of QuatE and OrthogonalE (3$\times$3) over different fixed entity dimensions $n*m$ on WN18RR, then Fig. 8 shows comparison of RotatE and OrthogonalE (2$\times$2) in FB15K-237. The experimental results, consistent with those discussed in the previous paragraph, demonstrate our model’s flexibility in controlling entity dimension through variable $m$ without unnecessarily increasing relation size.
@@
Furthermore, for OrthogonalE(2$\times$2) on WN18RR dataset, Fig. 5 result (with $m=7$ yielding MRR=0.483) suggests that a relation matrix of 40$\times$40 (20 parameters), compared to a dimension of 500$\times$500 (250 parameters) in Table 3, can achieve comparably high performance, thus demonstrating that entity matrix method significantly reduces the need for excessive relation parameters.
5.2.5 Relation Pattern
Following the proof of relation patterns in Appendix A.2, Fig. 9 shows histograms of relation embeddings for different relation patterns. We provide several relation patterns examples and discussion of non-commutative composition in Appendix A.3
Symmetry and Antisymmetry
In OrthogonalE, the symmetry relation pattern is encoded when the $\mathbf{e}_{R}$ embedding satisfies $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$, in accordance with Equation LABEL:equ:symmetery_proof. Figs. 9(a) and (b) illustrate the embeddings of $\mathbf{e}_{R}^{1}$ and $\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{R}^{1}-\mathrm{I}$, respectively. From Fig. 9(b), we observe that nearly all values are concentrated around 0, thereby indicating that OrthogonalE’s relations exhibit symmetry properties. Correspondingly, the multitude of non-zero values in Fig. 9(d) indicates that OrthogonalE’s relations also can represent antisymmetry properties.
Inversion
The inversion relation pattern is encoded when the $\mathbf{e}_{R}^{3}$ and $\mathbf{e}_{R}^{4}$ embedding satisfy $\mathbf{e}_{R}^{3}\cdot\mathbf{e}_{R}^{4}=\mathrm{I}$, in accordance with Equation LABEL:equ:inversion_proof. Even though $\mathbf{e}_{R}^{3}$ and $\mathbf{e}_{R}^{4}$ are responsible for additional relation patterns, which results in a cluster of values around -2 in Fig 9 (g), the majority of values still converge towards or equal 0. This suggests that OrthogonalE’s relations have the inversion property.
Composition
The composition relation pattern is encoded when the $\mathbf{e}_{R}^{5}$, $\mathbf{e}_{R}^{6}$, and $\mathbf{e}_{R}^{7}$ embedding satisfy $\mathbf{e}_{R}^{6}\cdot\mathbf{e}_{R}^{5}=\mathbf{e}_{R}^{7}$, in accordance with Equation LABEL:equ:composition_proof. The majority of data in Fig. 9 (k) converge towards or are equal to 0, indicating that OrthogonalE’s relations can represent the composition relation pattern.
@@ 6 Conclusion
In this study, we propose the OrthogonalE model to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Experimental results indicate that our new KGE model, OrthogonalE, offers generality and flexibility, captures several relation patterns, and significantly outperforms state-of-the-art KGE models while substantially reducing the number of relation parameters.
Limitations
Even though the Block-Diagonal Orthogonal relation with Riemannian Optimization makes KGE models more general and improves their performance, the computation of Exponential Retraction in the Orthogonal manifold for Riemannian Optimization is very costly. In practical model training, with the same entity dimension, our OrthogonalE (2$\times$2) training time is 4 times longer than that of RotatE. In future research directions, we will continue to explore this limitation, such as by employing the landing algorithm Ablin and Peyré (2022) for retraction on orthogonal manifolds to reduce computational complexity.
Ethics Statement
This study complies with the ACL Ethics Policy.
@@ Acknowledgements
Appendix A Appendix
Figure 10: Toy examples for symmetry, antisymmetry, inversion, and non-commutative composition relation patterns.
Figure 11: Histograms of relation embeddings for symmetry and antisymmetry relation patterns, where $\mathbf{e}_{R}^{1}$ represents _derivationally_related_form, $\mathbf{e}_{R}^{2}$ represents _instance_hypernym, $\mathbf{e}_{R}^{3}$ represents _also_see, $\mathbf{e}_{R}^{4}$ represents _verb_group, $\mathbf{e}_{R}^{5}$ represents /media_common/netflix_genre/titles, $\mathbf{e}_{R}^{6}$ represents /film/film/genre, $\mathbf{e}_{R}^{7}$ represents /award/award_category/category_of , and $\mathbf{e}_{R}^{8}$ represents /people/person/gender. From the WN18RR dataset, we select $\mathbf{e}_{R}^{1}$, $\mathbf{e}_{R}^{2}$ $\mathbf{e}_{R}^{3}$, $\mathbf{e}_{R}^{4}$and to represent Symmetry and Antisymmetry, respectively, and obtain their relation embeddings using the OrthogonalE(3$\times$3) model with $n$=501 and $m$=1. Similarly, from the FB15K-237 dataset, we select $\mathbf{e}_{R}^{5}$, $\mathbf{e}_{R}^{6}$, and $\mathbf{e}_{R}^{7}$, $\mathbf{e}_{R}^{8}$ as representations for symmetry and antisymmetr, respectively, and acquire their relation embeddings under the OrthogonalE(3$\times$3) model with $n$=999 and $m$=1.
Figure 12: Histograms of relation embeddings for composition relation patterns, where $\mathbf{e}_{R}^{1}$ represents /location/administrative_division/country, $\mathbf{e}_{R}^{2}$ represents /location/hud_county_place/place, $\mathbf{e}_{R}^{3}$ represents /base/aareas/schema/administrative_area/capital, $\mathbf{e}_{R}^{4}$ represents /award/award_nominee/award_nominations./award/award_nomination/nominated_for, $\mathbf{e}_{R}^{5}$ represents /award/award_category/winners./award/award_honor/award_winner, and $\mathbf{e}_{R}^{6}$ represents /award/award_category/nominees./award/award_nomination/nominated_for. $\mathbf{e}_{R}^{7}$, $\mathbf{e}_{R}^{8}$, and $\mathbf{e}_{R}^{9}$ have the same relational meanings as $\mathbf{e}_{R}^{4}$, $\mathbf{e}_{R}^{5}$, and $\mathbf{e}_{R}^{6}$, respectively, the difference lies in that the former are relations within the OrthogonalE(3$\times$3) model, while the latter are from the OrthogonalE(2$\times$2) model. All these relations are selected from the FB15K-237 dataset. $\mathbf{e}_{R}^{1}$, $\mathbf{e}_{R}^{2}$, $\mathbf{e}_{R}^{3}$, $\mathbf{e}_{R}^{7}$, $\mathbf{e}_{R}^{8}$, and $\mathbf{e}_{R}^{9}$ are relation embeddings under the OrthogonalE(3$\times$3) model with $n$=999 and $m$=1, while $\mathbf{e}_{R}^{4}$, $\mathbf{e}_{R}^{5}$, and $\mathbf{e}_{R}^{6}$ are relation embeddings under the OrthogonalE(2$\times$2) model with $n$=1000 and $m$=1
[…]
RotatE 0.1 - Adagrad 300
QuatE 0.2 - Adagrad 300
OrthogonalE(2$\times$2) 0.2 0.02 - 300
OrthogonalE(3$\times$3) 0.2 0.02 - 300
[…]
RotatE 0.1 - Adagrad 300
QuatE 0.2 - Adagrad 300
OrthogonalE(2$\times$2) 0.5 0.06 - 300
OrthogonalE(3$\times$3) 0.5 0.06 - 300
[…]
Table 5: Best hyperparameters of our approach and several composite models. In the table, the lr-entity values corresponding to TransE, RotatE, and QuatE refer to the learning rate for the entire model. For the OrthogonalE model, we employ RiemannianAdam for relation optimization and Adagrad for entity optimization, as detailed in the Approach section.
A.1 More information about Experiment setup
Dataset
WN18RR is a subset of WN18 Dettmers et al. (2018) which is contained in WordNet Miller (1995). FB15K-237 is a subset of FB15K, whichh is a subset of Freebase Bollacker et al. (2008), a comprehensive KG including data about common knowledge. All three datasets were designed for KGE, and we employ them for KGE tasks, and all three datasets have no individual people or offensive content.
Implementation
The training of models was carried out on two A6000 GPUs, which boasts 48GB of memory. Specifically, for the OrthogonalE model and its related flexible versions, the training durations were roughly 5 hour for the WN18RR dataset, 30 hours for FB15K-237. Our experiments were facilitated by leveraging PyTorch and Numpy as essential tools. Furthermore, We use ChatGPT in our paper writing. Finally, we obtain results by selecting the maximum values from three random seeds for Table 3 and using a single run for other results.
A.2 Proof of Relation Patterns
OrthogonalE is capable of representing the four kinds of relational patterns: Symmetry, Antisymmetry, Inversion, and Non-commutative Composition. We present the following propositions to formalize this capability:
Proposition 1.
OrthogonalE can represent Symmetry relation pattern.
Proof. If $(\mathbf{e}_{H},\mathbf{e}_{R},\mathbf{e}_{T})\in\mathcal{E}$ and $(\mathbf{e}_{T},\mathbf{e}_{R},\mathbf{e}_{H})\in\mathcal{E}$, we have
$\begin{split}&\mathbf{e}_{R}\cdot\mathbf{e}_{H}=\mathbf{e}_{T}\land\mathbf{e}_{R}\cdot\mathbf{e}_{T}=\mathbf{e}_{H}\\ &\Rightarrow\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}\\ &\Rightarrow\mathbf{X}_{i}\cdot\mathbf{X}_{i}=\mathrm{I}\\ \end{split}$ (8)
Proposition 2.
OrthogonalE can represent Antisymmetry relation pattern.
Proof. If $(\mathbf{e}_{H},\mathbf{e}_{R},\mathbf{e}_{T})\in\mathcal{E}$ and $(\mathbf{e}_{T},\mathbf{e}_{R},\mathbf{e}_{H})\notin\mathcal{E}$, we have
$\begin{split}&\mathbf{e}_{R}\cdot\mathbf{e}_{H}=\mathbf{e}_{T}\land\mathbf{e}_{R}\cdot\mathbf{e}_{T}\neq\mathbf{e}_{H}\\ &\Rightarrow\mathbf{e}_{R}\cdot\mathbf{e}_{R}\neq\mathrm{I}\\ &\Rightarrow\mathbf{X}_{i}\cdot\mathbf{X}_{i}\neq\mathrm{I}\\ \end{split}$ (9)
Proposition 3.
OrthogonalE can represent Inversion relation pattern.
Proof. If $(\mathbf{e}_{H},\mathbf{e}_{R}^{1},\mathbf{e}_{T})\in\mathcal{E}$ and $(\mathbf{e}_{T},\mathbf{e}_{R}^{2},\mathbf{e}_{H})\in\mathcal{E}$, we have
$\begin{split}&\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{H}=\mathbf{e}_{T}\land\mathbf{e}_{R}^{2}\cdot\mathbf{e}_{T}=\mathbf{e}_{H}\\ &\Rightarrow\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{R}^{2}=\mathrm{I}\\ &\Rightarrow\mathbf{X}_{i}^{1}\cdot\mathbf{X}_{i}^{2}=\mathrm{I}\\ \end{split}$ (10)
Proposition 4.
OrthogonalE can represent Non-commutative Composition relation pattern.
Proof. If $(\mathbf{e}_{H},\mathbf{e}_{R}^{1},\mathbf{e}_{T})\in\mathcal{E}$ and $(\mathbf{e}_{T},\mathbf{e}_{R}^{2},\mathbf{e}_{P})\in\mathcal{E}$ and $(\mathbf{e}_{H},\mathbf{e}_{R}^{3},\mathbf{e}_{P})\in\mathcal{E}$, we have
$\begin{split}&\mathbf{e}_{R}^{1}\!\cdot\!\mathbf{e}_{H}=\mathbf{e}_{T}\land\mathbf{e}_{R}^{2}\!\cdot\!\mathbf{e}_{T}=\mathbf{e}_{P}\land\mathbf{e}_{R}^{3}\!\cdot\!\mathbf{e}_{H}\!=\mathbf{e}_{P}\\ &\Rightarrow\mathbf{e}_{R}^{2}\cdot\mathbf{e}_{R}^{1}=\mathbf{e}_{R}^{3}\\ &\Rightarrow\mathbf{X}_{i}^{2}\cdot\mathbf{X}_{i}^{1}=\mathbf{X}_{i}^{3}\\ \end{split}$ (11)
$\mathbf{X}_{i}\in\mathbb{R}^{d\times d}\left\{\begin{array}[]{l}\text{is Commutative},\text{ if }d=2\\ \text{is Non-commutative},\text{ if }d>2\end{array}\right.$ (14)
The property of non-commutative composition dictates that the sequence of $\mathbf{X}_{i}^{1}$ and $\mathbf{X}_{i}^{2}$ cannot be exchanged. Given that $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ represents an orthogonal matrix, we consider two situations. In the first scenario, when $d=2$, $\mathbf{X}_{i}$ is a special case corresponding to a 2-dimensional rotation matrix, analogous to the RotatE Sun et al. (2019), and is therefore commutative, not exhibiting non-commutative composition. However, for $d>2$, $\mathbf{X}_{i}$ is non-commutative, thus can represent non-commutative composition relation pattern.
To gain a clearer understanding of the proof process, we use symmetry as an illustrative example to introduce the proof section, specifically referring to equation LABEL:equ:symmetery_proof in the paper. Initially, we assume that if a relation $\mathbf{e}_{R}$ in the OrthogonalE model exhibits the property of symmetry, then we can identify two related KG triples: $(\mathbf{e}_{H},\mathbf{e}_{R},\mathbf{e}_{T})\in\mathcal{E}$ and $(\mathbf{e}_{T},\mathbf{e}_{R},\mathbf{e}_{H})\in\mathcal{E}$. For instance, $\mathbf{e}_{H}$ ($\mathbf{e}_{R}$: is similar to) $\mathbf{e}_{T}$ and $\mathbf{e}_{T}$ ($\mathbf{e}_{R}$: is similar to) $\mathbf{e}_{H}$. Since both triples are trained by the OrthogonalE model, they adhere to the OrthogonalE equation (as depicted in Fig. 2). Consequently, we can derive that $\mathbf{e}_{R}\cdot\mathbf{e}_{H}=\mathbf{e}_{T}$ and $\mathbf{e}_{R}\cdot\mathbf{e}_{T}=\mathbf{e}_{H}$. By combining and simplifying these two equations, we can conclude that $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$ (Identity matrix). This means that if we can identify such an $\mathbf{e}_{R}$ that satisfies $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$, it demonstrates that the OrthogonalE model can capture the symmetry relation pattern. For $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$, we understand that $\mathbf{e}_{R}$ is composed of several block-diagonal orthogonal matrices $\mathbf{X}_{i}$, as shown in Fig. 2). Ultimately, this reduces to finding $\mathbf{X}_{i}\cdot\mathbf{X}_{i}=\mathrm{I}$ to satisfy $\mathbf{e}_{R}\cdot\mathbf{e}_{R}=\mathrm{I}$. We can identify corresponding orthogonal matrices $\mathbf{X}_{i}$ that satisfy $\mathbf{X}_{i}\cdot\mathbf{X}_{i}=\mathrm{I}$, which demonstrates that OrthogonalE can fulfill the property of symmetry.
A.3 More experiments on relation pattern
Symmetry and Antisymmetry
Fig. 11 shows histograms of additional examples of relation embeddings for symmetry and antisymmetry relation patterns. Furthermore, it displays examples of two symmetry and two antisymmetry relations from both the WN18RR and FB15K-237 datasets.
Composition
Firstly, we add $\mathbf{e}_{R}^{1}$, $\mathbf{e}_{R}^{2}$, and $\mathbf{e}_{R}^{3}$ from OrthogonalE(3$\times$3) for comparison with the three composition relations in Fig. 9 from OrthogonalE(2$\times$2). From Fig. 12(a, b, c, d), we observe that OrthogonalE(2$\times$2) performs better in composition than OrthogonalE(3$\times$3).
Secondly, we aim to explore more about the non-commutative composition relation pattern, so we select $\mathbf{e}_{R}^{4}$, $\mathbf{e}_{R}^{5}$, and $\mathbf{e}_{R}^{6}$, three non-commutative composition relations, for our study. Notably, $\mathbf{e}_{R}^{7}$, $\mathbf{e}_{R}^{8}$, and $\mathbf{e}_{R}^{9}$ share the same relational meanings as $\mathbf{e}_{R}^{4}$, $\mathbf{e}_{R}^{5}$, and $\mathbf{e}_{R}^{6}$, respectively, with the distinction that the former are relations within OrthogonalE(3$\times$3), while the latter are within OrthogonalE(2$\times$2). Figs. 12(h, l), using $\mathbf{e}_{R}^{5}\cdot\mathbf{e}_{R}^{4}-\mathbf{e}_{R}^{6}$ and $\mathbf{e}_{R}^{4}\cdot\mathbf{e}_{R}^{5}-\mathbf{e}_{R}^{6}$ respectively, show nearly indistinguishable histograms, indicating that swapping the sequence of the two relations in OrthogonalE(2$\times$2) does not affect the outcome of the composition, suggesting it is commutative. Conversely, Figs. 12(m, n), using $\mathbf{e}_{R}^{8}\cdot\mathbf{e}_{R}^{7}-\mathbf{e}_{R}^{9}$ and $\mathbf{e}_{R}^{7}\cdot\mathbf{e}_{R}^{8}-\mathbf{e}_{R}^{9}$, show that the former results in a trend closer to or equal to 0 more distinctly than the latter, implying that changing the sequence of relations affects the outcome, thereby demonstrating the non-commutative nature of relations in OrthogonalE(3$\times$3). In conclusion, even though OrthogonalE(2$\times$2) generally outperforms OrthogonalE(3$\times$3) in composition relation patterns, the comparative analysis reveals that OrthogonalE(3$\times$3) indeed possesses non-commutative composition properties, following the equation LABEL:equ:composition_proof and 14.
A.4 Introduction of Relation Patterns
We can observe several relation patterns in KGs, including symmetry, antisymmetry, inversion, and composition (both commutative and non-commutative). Detailed examples have been shown in Fig. 10.
Symmetry and Antisymmetry
Certain relations demonstrate symmetry, indicating that a relation’s validity between entities $x$ and $y$ ($\left(r_{1}(x,y)\Rightarrow r_{1}(y,x)\right)$) (for instance, is married to) is equally valid in the opposite direction (namely, from $y$ to $x$). Conversely, other relations are characterized by antisymmetry ($\left(r_{1}(x,y)\Rightarrow\neg r_{1}(y,x)\right)$), signifying that if a relation is applicable between $x$ and $y$ (such as is father of), it is inapplicable in the reverse direction (from $y$ to $x$).
Inversion
Relations can also exhibit inversion ($\left(r_{1}(x,y)\Leftrightarrow r_{2}(y,x)\right)$), where reversing the direction of a relation effectively transforms it into another relation (for example, is child of and is parent of).
Composition
The composition of relations ($\left(r_{1}(x,y)\cap r_{2}(y,z)\Rightarrow r_{3}(x,z)\right)$) signifies a crucial pattern wherein merging two or more relations facilitates the deduction of a novel relation. Such compositions might be commutative, where the sequence of relations is irrelevant, or non-commutative, where the sequence significantly influences the outcome. In scenarios where the order of relations is pivotal, as illustrated by the relationship where B is the mother of A’s father and E is the father of A’s mother, non-commutative composition ($\left(r_{1}(x,y)\cap r_{2}(y,z)\neq\right(r_{2}(x,y)\cap r_{1}(y,z)$) becomes essential. While commutative compositions would consider B and E as identical, non-commutative compositions recognize them as distinct.
[…]
SimKGCWang et al. (2022a) Language Models .671
LERPHan et al. (2023) Additional Contextual Information (Logic Rules) .622
C-LMKEWang et al. (2022b) Language Models .598
KNN-KGEZhang et al. (2022) Language Models .579
HittERChen et al. (2020) Language Models .503
OrthogonalE($10\times 10$) - .494
[…]
Table 6: Other baseline models in WN18RR dataset.
A.5 Other baseline KGE model
In recent times, several significant performance methods have been developed, as detailed for WN18RR in Table 6. Among these, MoCoSAHe et al. (2023), SimKGCWang et al. (2022a), C-LMKEWang et al. (2022b), KNN-KGEZhang et al. (2022), and HittERChen et al. (2020) mainly utilize Language Models (LMs) to enrich dataset semantic information, thereby achieving superior outcomes. Conversely, LERPHan et al. (2023) does not employ LMs but leverages additional contextual information (logic rules) beyond the dataset to fill information gaps in entities and relations. On the other hand, methods such as TransEBordes et al. (2013), RotatESun et al. (2019), and the OrthogonalE method introduced in this paper depend solely on the inherent data and information of the KGE dataset itself. These methods, based on specific mathematical rules and algorithms, do not incorporate any external information and thus do not operate as black-box approaches like LLMs. Consequently, these dataset-dependent methods remain highly valuable for KGE research.
A.6 hyperparameter
All the hyperparameter settings have been shown in Table 5